OpenStudy (anonymous):

What woulf 3g+5=17 what eould g be

OpenStudy (michele_laino):

we can add -5 to both sides, so we can write: \[3g+5-5=17-5\] then \[3g=12\] now divide by 3 both sides, please!

OpenStudy (anonymous):

First, you subtract 5 from both sides. 3g = 12 Then, you divide both sides by 3. g = 4 Therefore, the g is equal to 4